# Dyson Equation for Correlated Linearizations and Test Error of Random Features Regression

### v2

@@ 6.1 First Deterministic Equivalent
$(2(\delta-\Re[z])^{-2}d^{2}\|K_{AA^{T}}\|^{2}-1-\Re[z])\epsilon\leq 2^{-1}(1+\Re[z])$
and assume, by contradiction, that $d\|K_{AA^{T}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{T}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}>1-\epsilon$. Using the definition of $M_{1,1}$ and $M_{2,2}$ and repeatedly applying (24),
$\mathrm{tr}(K_{AA^{T}}\Re[M_{1,1}])$ $=(\delta-\Re[z])\mathrm{tr}(K_{AA^{T}}M_{1,1}M_{1,1}^{\ast})$
$-\mathrm{tr}(\Re[M_{2,2}])\mathrm{tr}(K_{AA^{T}}M_{1,1}K_{AA^{T}}M_{1,1}^{\ast})$
$=(\delta-\Re[z])\mathrm{tr}(K_{AA^{T}}M_{1,1}M_{1,1}^{\ast})$
$+d\|M_{2,2}\|^{2}\|K_{AA^{T}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{T}}^{\frac{1}{2}}\|^{2}_{F}(1+\Re[z])$
$+d\|M_{2,2}\|^{2}\|K_{AA^{T}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{T}}^{\frac{1}{2}}\|^{2}_{F}\mathrm{tr}(K_{AA^{T}}\Re[M_{1,1}])$
$\geq(\delta-\Re[z])\mathrm{tr}(K_{AA^{T}}M_{1,1}M_{1,1}^{\ast})+(1-\epsilon)(1+\Re[z])$
$+(1-\epsilon)\mathrm{tr}(K_{AA^{T}}\Re[M_{1,1}]).$
Solving for $\mathrm{tr}(K_{AA^{T}}\Re[M_{1,1}])$, we obtain that $\mathrm{tr}(K_{AA^{T}}\Re[M_{1,1}])=t\epsilon^{-1}$ with
$t:=(\delta-\Re[z])\mathrm{tr}(K_{AA^{T}}M_{1,1}M_{1,1}^{\ast})+(1-\epsilon)(1+\Re[z]).$
@@
$-\Re[M_{2,2}]=\|M_{2,2}\|^{2}(1+\Re[z]+\mathrm{tr}(K_{AA^{T}}\Re[M_{1,1}]))I_{d}\succeq\|M_{2,2}\|^{2}(1+\Re[z]+t\epsilon^{-1})I_{d}.$
@@
$d\|K_{AA^{T}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{T}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}$ $\leq\frac{(\delta-\Re[z])^{-2}d^{2}\|K_{AA^{T}}\|^{2}}{1+\Re[z]+t\epsilon^{-1}}\leq 2^{-1}.$
@@
$1-d\|K_{AA^{T}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{T}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}\geq\epsilon$
@@
The statement of Lemma 6.7 is intricate because the left-hand side of the inequality $(2\delta^{-2}d^{2}\|K_{AA^{T}}\|^{2}-1-\Re[z])\epsilon\leq 2^{-1}(1+\Re[z])$ may be negative. Nevertheless, the essence of Lemma 6.7 lies in the fact that the quantity $1-d\|K_{AA^{T}}^{1/2}M_{1,1}(z)K_{AA^{T}}^{1/2}\|^{2}_{F}\|M_{2,2}(z)\|^{2}$ can be consistently bounded away from 00 regardless of the dimension.
Now that we have some control on the solution of the DEL when the spectral parameter is close to the origin, we still need to continuously extend the function $M$ to its boundary point 00. To do so, we analytically extend $M$ by reflection to the lower complex plane $\{z\in\mathbb{H}\,:\,\Im[z]<0\}$ through an open interval containing the origin.
@@
Using the definition of matrix imaginary part and the resolvent identity, we obtain the system of equations $\Im[M_{1,1}]=M_{1,1}(\Im[z]+\mathrm{tr}(\Im[M_{2,2}])K_{AA^{T}})(M_{1,1})^{\ast}$ and $\mathrm{tr}(\Im[M_{2,2}])=d\|M_{2,2}\|^{2}(\Im[z]+\mathrm{tr}(K_{AA^{T}}\Im[M_{1,1}]))$. Combining the two equalities, we get $d^{-1}\mathrm{tr}(\Im[M_{2,2}])(1-\|\sqrt{d}K_{AA^{T}}^{1/2}M_{1,1}K_{AA^{T}}^{1/2}\|^{2}_{F}\|M_{2,2}\|^{2})=\|M_{2,2}\|^{2}\Im[z](1+\|\sqrt{d}K_{AA^{T}}^{1/2}M_{1,1}\|_{F}^{2})$. By Lemma 6.7, $01-d\|K_{AA^{T}}^{\frac{1}{2}}M_{1,1}K_{AA^{T}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}\|^{2}>$ 0 uniformly on $\{z\in\mathbb{H}\,:\,|z|\leq\epsilon\}$ for every $0<\epsilon<1\wedge\delta$. Using Lemma 6.3,
$d^{-1}\mathrm{tr}(\Im[M_{2,2}])\leq\frac{\Im[z](1+\|\sqrt{d}(\delta-\Re[z])^{-1}K_{AA^{T}}^{\frac{1}{2}}\|_{F}^{2})}{1-\|\sqrt{d}K_{AA^{T}}^{\frac{1}{2}}M_{1,1}K_{AA^{T}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}\|^{2}}.$
Thus, we observe that $\Im[M_{2,2}(z)]\downarrow 0$ uniformly as $\Im[z]\to 0$ on $(-\epsilon,\epsilon)$ and similarly for $\|\Im[M_{1,1}]\|$. Since $M_{1,1}$ and $M_{2,2}$ fully define the solution of the DEL, $\|\Im[M(z)]\|$ vanishes uniformly for $\Re[z]\in(-\epsilon,\epsilon)$ as $\Im[z]\to 0$. By the Stieltjes inversion lemma [30, Theorem 5.4(v) and Theorem 5.4(vi)], the positive semidefinite measure in Theorem 2.1 has no support in $(-\epsilon,\epsilon)$. The result follows from [30, Lemma 5.6]. ∎
@@
$T:x\in\mathbb{R}_{<0}\mapsto-(1+\mathrm{tr}(K_{AA^{T}}(\delta I_{n_{\text{train}}}-dxK_{AA^{T}})^{-1}))^{-1}\in\mathbb{R}_{<0}$
@@
$M(0)=\begin{bmatrix}(\delta I_{n_{\text{train}}}-d\alpha K_{AA^{T}})^{-1}&0&-d\alpha M_{1,1}(0)K_{A\hat{A}^{T}}&0\\ 0&\alpha I_{d}&0&0\\ -d\alpha K_{\hat{A}A^{T}}M_{1,1}(0)&0&(d\alpha)^{2}K_{\hat{A}A^{T}}M_{1,1}(0)K_{A\hat{A}^{T}}+d\alpha K_{\hat{A}\hat{A}^{T}}&-I_{n_{\text{test}}}\\ 0&0&-I_{n_{\text{test}}}&0\end{bmatrix}$
@@
In order to use the Earle-Hamilton fixed-point theorem [23], we consider the set $\mathcal{S}$ of complex matrices $N\in\mathbb{C}^{(n_{\text{train}}+d)\times(n_{\text{train}}+d)}$ with $N_{1,2}=N_{2,1}=0$, $\Re[N_{1,1}]\succ 0$ and $\Re[N_{2,2}]\prec 0$ and denote
@@
Using an argument analogous to the one in lemmas 6.3 and 4.14, we see that
$\Re[\mathcal{F}^{(\mathrm{sub})}_{1,1}(N)]\succeq\delta\mathcal{F}^{(\mathrm{sub})}_{1,1}(N)(\mathcal{F}^{(\mathrm{sub})}_{1,1}(N))^{\ast}\succeq\delta(\delta+d\|K_{AA^{T}}\|\|N\|)^{-2}$
@@
$\Re[\mathcal{F}^{(\mathrm{sub})}_{2,2}(N)]\preceq-\mathcal{F}^{(\mathrm{sub})}_{2,2}(N)(\mathcal{F}^{(\mathrm{sub})}_{2,2}(N))^{\ast}\preceq-(1+d\|K_{AA^{T}}\|\|N\|)^{-2}$
@@ 6.2 Second Deterministic Equivalent
We now consider the squared matrix $(AA^{T}+\delta I_{n_{\text{train}}})^{-1}A\hat{A}^{T}\hat{A}A^{T}(AA^{T}+\delta I_{n_{\text{train}}})^{-1}$. Notice that $(L^{-2})_{1,1}=R^{2}+RA^{T}AR+RA\hat{A}^{T}\hat{A}AR$ and $(L^{(\mathrm{sub})})^{-2}_{1,1}=R^{2}+RA^{T}AR$ for $R:=(AA^{T}+\delta I_{n_{\text{train}}})^{-1}$. Rearranging, we get that
$(AA^{T}+\delta I_{n_{\text{train}}})^{-1}A\hat{A}^{T}\hat{A}A(AA^{T}+\delta I_{n_{\text{train}}})^{-1}=(L^{-2})_{1,1}-(L^{(\mathrm{sub})})^{-2}_{1,1}.$ (25)
Therefore, it suffices to find deterministic equivalents for $(L^{(\mathrm{sub})})^{-2}$ and $L^{-2}$ to obtain a deterministic equivalent for the random matrix $(AA^{T}+\delta I_{n_{\text{train}}})^{-1}A\hat{A}^{T}\hat{A}A(AA^{T}+\delta I_{n_{\text{train}}})^{-1}$.
@@
Under the settings of Theorem 2.4, let $\alpha=d^{-1}\mathrm{tr}(M_{2,2}(0))$ as in Lemma 6.10, $R=(AA^{T}+\delta I_{n_{\text{train}}})^{-1}$ and define
$\beta=\frac{\alpha^{2}\mathrm{tr}\left(K_{\hat{A}\hat{A}^{T}}+d\alpha K_{\hat{A}A^{T}}M_{1,1}(0)(I_{n_{\text{train}}}+\delta M_{1,1}(0))K_{A\hat{A}^{T}}\right)}{1-\|\sqrt{d}\alpha K_{AA^{T}}^{\frac{1}{2}}M_{1,1}(0)K_{AA^{T}}^{\frac{1}{2}}\|_{F}^{2}}\in\mathbb{R}_{\geq 0}.$
Then, $\mathrm{tr}U(RA\hat{A}^{T}\hat{A}A^{T}R-d\beta M_{1,1}(0)K_{AA^{T}}M_{1,1}(0)-M_{1,3}(0)M_{3,1}(0))\to 0$ almost surely as $n\to\infty$ for every sequence $U\in\mathbb{C}^{n_{\text{train}}\times n_{\text{train}}}$ with $\|U\|_{\ast}\leq 1$.
@@
First, as stated in Lemma 4.16, we note that $i\tau\mapsto(L-i\tau)^{-1}$ is an analytic function with $\partial_{i\tau}(L-i\tau)^{-1}=(L-i\tau)^{-2}$. Overloading notation, let $M^{(\zeta)}\in\mathcal{M}_{+}$ be the unique solution to the DEL $(\mathbb{E}L-\mathcal{S}(M^{(\zeta)})-\zeta I_{\ell})M^{(\zeta)}=I_{\ell}$ where $\zeta\in\mathbb{H}$. By the proof of [26, Theorem 2.14], the function $\zeta\mapsto M^{(\zeta)}$ is analytic on $\mathbb{H}$. Adapting a general argument resembling to the one in [56, equation (174)], it follows from Cauchy’s integral formula that
@@
where $\gamma$ forms a counterclockwise circle of radius $\tau/2$ around $i\tau$. We know that $M^{(\zeta)}$ is a deterministic equivalent for $(L-\zeta)^{-1}$ for every fixed $\zeta\in\mathbb{H}$. By the resolvent identity, $\zeta\mapsto(L-\zeta I_{\ell})^{-1}$ is $4/\tau^{2}$-Lipschitz on $\{z\in\mathbb{H}\,:\,\Im[z]\geq\tau/2\}$. Similarly, by the proof of [26, Theorem 2.14], the function $\zeta\mapsto M^{(\zeta)}$ is $(2/\tau)^{12}$-Lipschitz on $\{z\in\mathbb{H}\,:\,\Im[z]\geq\tau/2\}$. Therefore, we obtain $\mathrm{tr}(U((L-i\tau)^{-2}-\partial_{i\tau}M^{(\tau)}(0)))\to 0$ almost surely as $n\to\infty$ for every $\tau\in\mathbb{R}_{>0}$ and $U\in\mathbb{C}^{\ell\times\ell}$ with $\|U\|_{\ast}\leq 1$. Taking the derivative of (5), we obtain $\partial_{i\tau}M^{(\tau)}(0)=M^{(\tau)}(0)(\mathcal{S}(\partial_{i\tau}M^{(\tau)}(0))+I_{\ell})M^{(\tau)}(0)$ or, relating this equation to the stability operator, $\mathcal{L}^{(\tau)}(\partial_{i\tau}M^{(\tau)}(0))=(M^{(\tau)}(0))^{2}$ with $\mathcal{L}^{(\tau)}:N\in\mathbb{C}^{\ell\times\ell}\mapsto N-M^{(\tau)}(0)\mathcal{S}(N)M^{(\tau)}(0)$.
@@
$=\mathrm{tr}((M_{2,2}^{(\tau)})^{2})(\mathrm{tr}(K_{AA^{T}}C_{1,1}+K_{A\hat{A}^{T}}C_{1,4}^{T}+K_{\hat{A}A^{T}}C_{1,4}+K_{\hat{A}\hat{A}^{T}}C_{4,4})+1)$
$+\mathrm{tr}(\partial_{i\tau}M_{2,2}^{(\tau)})\mathrm{tr}((M_{2,2}^{(\tau)})^{2})\mathrm{tr}(K_{AA^{T}}D_{1,1}+K_{A\hat{A}^{T}}D_{1,4}^{T}+K_{\hat{A}A^{T}}D_{1,4}+K_{\hat{A}\hat{A}^{T}}D_{4,4}).$
By the proof of Lemma 6.6, we observe that there exists a function $f:\mathbb{R}_{>0}\mapsto\mathbb{R}_{\geq 0}$ with $\lim_{\tau\downarrow 0}f(\tau)=0$ such that $\|D_{4,4}\|\leq f(\tau)+o_{n}(1)$, $\|D_{1,1}-M_{1,1}K_{AA^{T}}M_{1,1}\|\leq f(\tau)+o_{n}(1)$, $\|D_{1,4}\|=\|D_{4,1}^{T}\|\leq f(\tau)+o_{n}(1)$ and $\|M_{2,2}^{(\tau)}-M_{2,2}\|\leq f(\tau)+o_{n}(1)$. By Lemma 6.7,
$|1-\mathrm{tr}((M_{2,2}^{(\tau)})^{2})\mathrm{tr}(K_{AA^{T}}D_{1,1}+K_{A\hat{A}^{T}}D_{4,1}+K_{\hat{A}A^{T}}D_{1,4}+K_{\hat{A}\hat{A}^{T}}D_{4,4})|$
is bounded away from 00 for every $n\in\mathbb{N}$ large enough and $\tau\in\mathbb{R}_{>0}$ small enough. In particular, the limit $\lim_{\tau\to 0}\partial_{i\tau}M^{(\tau)}$ exists and satisfies
$\lim_{\tau\to 0}d^{-1}\mathrm{tr}(\partial_{i\tau}M_{2,2}^{(\tau)})$ $=\frac{\alpha^{2}(\mathrm{tr}(K_{AA^{T}}M^{2}_{1,1}+K_{AA^{T}}M_{1,3}M_{3,1}+K_{\hat{A}\hat{A}^{T}})+1)}{1-\|\sqrt{d}\alpha K_{AA^{T}}^{\frac{1}{2}}M_{1,1}K_{AA^{T}}^{\frac{1}{2}}\|_{F}^{2}}$
$-\frac{\alpha^{2}\mathrm{tr}(K_{A\hat{A}^{T}}M_{3,1}+K_{\hat{A}A^{T}}M_{1,3})}{1-\|\sqrt{d}\alpha K_{AA^{T}}^{\frac{1}{2}}M_{1,1}K_{AA^{T}}^{\frac{1}{2}}\|_{F}^{2}}$
$=\beta+\frac{\alpha^{2}(1+\mathrm{tr}(K_{AA^{T}}M^{2}_{1,1}))}{1-\|\sqrt{d}\alpha K_{AA^{T}}^{\frac{1}{2}}M_{1,1}K_{AA^{T}}^{\frac{1}{2}}\|_{F}^{2}}$

### v4

@@ 6.1 First Deterministic Equivalent
$(2(\delta-\Re[z])^{-2}d^{2}\|K_{AA^{\top}}\|^{2}-1-\Re[z])\epsilon\leq 2^{-1}(1+\Re[z])$
and assume, by contradiction, that $d\|K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{\top}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}>1-\epsilon$. Using the definition of $M_{1,1}$ and $M_{2,2}$ and repeatedly applying (24),
$\mathrm{tr}(K_{AA^{\top}}\Re[M_{1,1}])$ $=(\delta-\Re[z])\mathrm{tr}(K_{AA^{\top}}M_{1,1}M_{1,1}^{\ast})$
$-\mathrm{tr}(\Re[M_{2,2}])\mathrm{tr}(K_{AA^{\top}}M_{1,1}K_{AA^{\top}}M_{1,1}^{\ast})$
$=(\delta-\Re[z])\mathrm{tr}(K_{AA^{\top}}M_{1,1}M_{1,1}^{\ast})$
$+d\|M_{2,2}\|^{2}\|K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{\top}}^{\frac{1}{2}}\|^{2}_{F}(1+\Re[z])$
$+d\|M_{2,2}\|^{2}\|K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{\top}}^{\frac{1}{2}}\|^{2}_{F}\mathrm{tr}(K_{AA^{\top}}\Re[M_{1,1}])$
$\geq(\delta-\Re[z])\mathrm{tr}(K_{AA^{\top}}M_{1,1}M_{1,1}^{\ast})+(1-\epsilon)(1+\Re[z])$
$+(1-\epsilon)\mathrm{tr}(K_{AA^{\top}}\Re[M_{1,1}]).$
Solving for $\mathrm{tr}(K_{AA^{\top}}\Re[M_{1,1}])$, we obtain that $\mathrm{tr}(K_{AA^{\top}}\Re[M_{1,1}])=t\epsilon^{-1}$ with
$t:=(\delta-\Re[z])\mathrm{tr}(K_{AA^{\top}}M_{1,1}M_{1,1}^{\ast})+(1-\epsilon)(1+\Re[z]).$
@@
$\Re[M_{2,2}]=\|M_{2,2}\|^{2}(1+\Re[z]+\mathrm{tr}(K_{AA^{\top}}\Re[M_{1,1}]))I_{d}\succeq\|M_{2,2}\|^{2}(1+\Re[z]+t\epsilon^{-1})I_{d}.$
@@
$d\|K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{\top}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}$ $\leq\frac{(\delta-\Re[z])^{-2}d^{2}\|K_{AA^{\top}}\|^{2}}{1+\Re[z]+t\epsilon^{-1}}\leq 2^{-1}.$
@@
$1-d\|K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}(z)K_{AA^{\top}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}(z)\|^{2}\geq\epsilon$
@@
The statement of Lemma 6.7 is intricate because the left-hand side of the inequality
$(2\delta^{-2}d^{2}\|K_{AA^{\top}}\|^{2}-1-\Re[z])\epsilon\leq 2^{-1}(1+\Re[z])$
may be negative. Nevertheless, the essence of Lemma 6.7 lies in the fact that the quantity $1-d\|K_{AA^{\top}}^{1/2}M_{1,1}(z)K_{AA^{\top}}^{1/2}\|^{2}_{F}\|M_{2,2}(z)\|^{2}$ can be consistently bounded away from 0 regardless of the dimension.
Now that we have some control on the solution of the DEL when the spectral parameter is close to the origin, we still need to continuously extend the function $M$ to its boundary point 0. To do so, we analytically extend $M$ by reflection to the lower complex plane $\{z\in\mathbb{H}\,:\,\Im[z]<0\}$ through an open interval containing the origin.
@@
Using the definition of matrix imaginary part and the resolvent identity, we obtain the system of equations $\Im[M_{1,1}]=M_{1,1}(\Im[z]+\mathrm{tr}(\Im[M_{2,2}])K_{AA^{\top}})(M_{1,1})^{\ast}$ and $\mathrm{tr}(\Im[M_{2,2}])=d\|M_{2,2}\|^{2}(\Im[z]+\mathrm{tr}(K_{AA^{\top}}\Im[M_{1,1}]))$. Combining the two equalities, we get $d^{-1}\mathrm{tr}(\Im[M_{2,2}])(1-\|\sqrt{d}K_{AA^{\top}}^{1/2}M_{1,1}K_{AA^{\top}}^{1/2}\|^{2}_{F}\|M_{2,2}\|^{2})=\|M_{2,2}\|^{2}\Im[z](1+\|\sqrt{d}K_{AA^{\top}}^{1/2}M_{1,1}\|_{F}^{2})$. By Lemma 6.7, $1-d\|K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}K_{AA^{\top}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}\|^{2}>0$ uniformly on $\{z\in\mathbb{H}\,:\,|z|\leq\epsilon\}$ for every $0<\epsilon<1\wedge\delta$. Using Lemma 6.3,
$d^{-1}\mathrm{tr}(\Im[M_{2,2}])\leq\frac{\Im[z](1+\|\sqrt{d}(\delta-\Re[z])^{-1}K_{AA^{\top}}^{\frac{1}{2}}\|_{F}^{2})}{1-\|\sqrt{d}K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}K_{AA^{\top}}^{\frac{1}{2}}\|^{2}_{F}\|M_{2,2}\|^{2}}.$
Thus, we observe that $\Im[M_{2,2}(z)]\downarrow 0$ uniformly as $\Im[z]\to 0$ on $(-\epsilon,\epsilon)$ and similarly for $\|\Im[M_{1,1}]\|$. Since $M_{1,1}$ and $M_{2,2}$ fully define the solution of the DEL, $\|\Im[M(z)]\|$ vanishes uniformly for $\Re[z]\in(-\epsilon,\epsilon)$ as $\Im[z]\to 0$. By the Stieltjes inversion lemma [matrix_herglotz, Theorem 5.4(v) and Theorem 5.4(vi)], the positive semidefinite measure in Theorem 2.1 has no support in $(-\epsilon,\epsilon)$. The result follows from [matrix_herglotz, Lemma 5.6]. ∎
@@
$T:x\in\mathbb{R}_{<0}\mapsto-(1+\mathrm{tr}(K_{AA^{\top}}(\delta I_{n_{\text{train}}}-dxK_{AA^{\top}})^{-1}))^{-1}\in\mathbb{R}_{<0}$
@@
$M(0)=\begin{bmatrix}(\delta I_{n_{\text{train}}}-d\alpha K_{AA^{\top}})^{-1}&0&-d\alpha M_{1,1}(0)K_{A\hat{A}^{\top}}&0\\ 0&\alpha I_{d}&0&0\\ -d\alpha K_{\hat{A}A^{\top}}M_{1,1}(0)&0&(d\alpha)^{2}K_{\hat{A}A^{\top}}M_{1,1}(0)K_{A\hat{A}^{\top}}+d\alpha K_{\hat{A}\hat{A}^{\top}}&-I_{n_{\text{test}}}\\ 0&0&-I_{n_{\text{test}}}&0\end{bmatrix}$
@@
In order to use the Earle-Hamilton fixed-point theorem [earl_hamilton], we consider the set $\mathcal{S}$ of complex matrices $N\in\mathbb{C}^{(n_{\text{train}}+d)\times(n_{\text{train}}+d)}$ with $N_{1,2}=N_{2,1}=0$, $\Re[N_{1,1}]\succ 0$ and $\Re[N_{2,2}]\prec 0$ and denote
@@
Using an argument analogous to the one in lemmas 6.3 and 4.10, we see that
$\Re[\mathcal{F}^{(\mathrm{sub})}_{1,1}(N)]\succeq\delta\mathcal{F}^{(\mathrm{sub})}_{1,1}(N)(\mathcal{F}^{(\mathrm{sub})}_{1,1}(N))^{\ast}\succeq\delta(\delta+d\|K_{AA^{\top}}\|\|N\|)^{-2}$
@@
$\Re[\mathcal{F}^{(\mathrm{sub})}_{2,2}(N)]\preceq-\mathcal{F}^{(\mathrm{sub})}_{2,2}(N)(\mathcal{F}^{(\mathrm{sub})}_{2,2}(N))^{\ast}\preceq-(1+d\|K_{AA^{\top}}\|\|N\|)^{-2}$
@@ 6.2 Second Deterministic Equivalent
We now consider the squared matrix $(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}A\hat{A}^{\top}\hat{A}A^{\top}(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}$. Notice that $(L^{-2})_{1,1}=R^{2}+RA^{\top}AR+RA\hat{A}^{\top}\hat{A}AR$ and $(L^{(\mathrm{sub})})^{-2}_{1,1}=R^{2}+RA^{\top}AR$ for $R:=(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}$. Rearranging, we get that
$(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}A\hat{A}^{\top}\hat{A}A(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}=(L^{-2})_{1,1}-(L^{(\mathrm{sub})})^{-2}_{1,1}.$ (25)
Therefore, it suffices to find deterministic equivalents for $(L^{(\mathrm{sub})})^{-2}$ and $L^{-2}$ to obtain a deterministic equivalent for the random matrix $(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}A\hat{A}^{\top}\hat{A}A(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}$.
@@
Under the settings of Theorem 2.3, let $\alpha=d^{-1}\mathrm{tr}(M_{2,2}(0))$ as in Lemma 6.9, $R=(AA^{\top}+\delta I_{n_{\text{train}}})^{-1}$ and define
$\beta=\frac{\alpha^{2}\mathrm{tr}\left(K_{\hat{A}\hat{A}^{\top}}+d\alpha K_{\hat{A}A^{\top}}M_{1,1}(0)(I_{n_{\text{train}}}+\delta M_{1,1}(0))K_{A\hat{A}^{\top}}\right)}{1-\|\sqrt{d}\alpha K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}(0)K_{AA^{\top}}^{\frac{1}{2}}\|_{F}^{2}}\in\mathbb{R}_{\geq 0}.$
Then, $\mathrm{tr}U(RA\hat{A}^{\top}\hat{A}A^{\top}R-d\beta M_{1,1}(0)K_{AA^{\top}}M_{1,1}(0)-M_{1,3}(0)M_{3,1}(0))\to 0$ almost surely as $n\to\infty$ for every sequence $U\in\mathbb{C}^{n_{\text{train}}\times n_{\text{train}}}$ with $\|U\|_{\ast}\leq 1$.
@@
First, as stated in Lemma 4.11, we note that $i\tau\mapsto(L-i\tau)^{-1}$ is an analytic function with $\partial_{i\tau}(L-i\tau)^{-1}=(L-i\tau)^{-2}$. Overloading notation, let $M^{(\zeta)}\in\mathcal{M}_{+}$ be the unique solution to the DEL $(\mathbb{E}L-\mathcal{S}(M^{(\zeta)})-\zeta I_{\ell})M^{(\zeta)}=I_{\ell}$ where $\zeta\in\mathbb{H}$. By the proof of [nemish_local_2020, Theorem 2.14], the function $\zeta\mapsto M^{(\zeta)}$ is analytic on $\mathbb{H}$. Adapting a general argument resembling to the one in [schroder2023deterministic, equation (174)], it follows from Cauchy’s integral formula that
@@
where $\gamma$ forms a counterclockwise circle of radius $\tau/2$ around $i\tau$. We know that $M^{(\zeta)}$ is a deterministic equivalent for $(L-\zeta)^{-1}$ for every fixed $\zeta\in\mathbb{H}$. By the resolvent identity, $\zeta\mapsto(L-\zeta I_{\ell})^{-1}$ is $4/\tau^{2}$-Lipschitz on $\{z\in\mathbb{H}\,:\,\Im[z]\geq\tau/2\}$. Similarly, by the proof of [nemish_local_2020, Theorem 2.14], the function $\zeta\mapsto M^{(\zeta)}$ is $(2/\tau)^{12}$-Lipschitz on $\{z\in\mathbb{H}\,:\,\Im[z]\geq\tau/2\}$. Therefore, we obtain $\mathrm{tr}(U((L-i\tau)^{-2}-\partial_{i\tau}M^{(\tau)}(0)))\to 0$ almost surely as $n\to\infty$ for every $\tau\in\mathbb{R}_{>0}$ and $U\in\mathbb{C}^{\ell\times\ell}$ with $\|U\|_{\ast}\leq 1$. Taking the derivative of (5), we obtain $\partial_{i\tau}M^{(\tau)}(0)=M^{(\tau)}(0)(\mathcal{S}(\partial_{i\tau}M^{(\tau)}(0))+I_{\ell})M^{(\tau)}(0)$ or, relating this equation to the stability operator, $\mathcal{L}^{(\tau)}(\partial_{i\tau}M^{(\tau)}(0))=(M^{(\tau)}(0))^{2}$ with $\mathcal{L}^{(\tau)}:N\in\mathbb{C}^{\ell\times\ell}\mapsto N-M^{(\tau)}(0)\mathcal{S}(N)M^{(\tau)}(0)$.
@@
$=\mathrm{tr}((M_{2,2}^{(\tau)})^{2})(\mathrm{tr}(K_{AA^{\top}}C_{1,1}+K_{A\hat{A}^{\top}}C_{1,4}^{\top}+K_{\hat{A}A^{\top}}C_{1,4}+K_{\hat{A}\hat{A}^{\top}}C_{4,4})+1)$
$+\mathrm{tr}(\partial_{i\tau}M_{2,2}^{(\tau)})\mathrm{tr}((M_{2,2}^{(\tau)})^{2})\mathrm{tr}(K_{AA^{\top}}D_{1,1}+K_{A\hat{A}^{\top}}D_{1,4}^{\top}+K_{\hat{A}A^{\top}}D_{1,4}+K_{\hat{A}\hat{A}^{\top}}D_{4,4}).$
By the proof of Lemma 6.6, we observe that there exists a function $f:\mathbb{R}_{>0}\mapsto\mathbb{R}_{\geq 0}$ with $\lim_{\tau\downarrow 0}f(\tau)=0$ such that $\|D_{4,4}\|\leq f(\tau)+o_{n}(1)$, $\|D_{1,1}-M_{1,1}K_{AA^{\top}}M_{1,1}\|\leq f(\tau)+o_{n}(1)$, $\|D_{1,4}\|=\|D_{4,1}^{\top}\|\leq f(\tau)+o_{n}(1)$ and $\|M_{2,2}^{(\tau)}-M_{2,2}\|\leq f(\tau)+o_{n}(1)$. By Lemma 6.7,
$|1-\mathrm{tr}((M_{2,2}^{(\tau)})^{2})\mathrm{tr}(K_{AA^{\top}}D_{1,1}+K_{A\hat{A}^{\top}}D_{4,1}+K_{\hat{A}A^{\top}}D_{1,4}+K_{\hat{A}\hat{A}^{\top}}D_{4,4})|$
is bounded away from 0 for every $n\in\mathbb{N}$ large enough and $\tau\in\mathbb{R}_{>0}$ small enough. In particular, the limit $\lim_{\tau\to 0}\partial_{i\tau}M^{(\tau)}$ exists and satisfies
$\lim_{\tau\to 0}d^{-1}\mathrm{tr}(\partial_{i\tau}M_{2,2}^{(\tau)})$ $=\frac{\alpha^{2}(\mathrm{tr}(K_{AA^{\top}}M^{2}_{1,1}+K_{AA^{\top}}M_{1,3}M_{3,1}+K_{\hat{A}\hat{A}^{\top}})+1)}{1-\|\sqrt{d}\alpha K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}K_{AA^{\top}}^{\frac{1}{2}}\|_{F}^{2}}$
$-\frac{\alpha^{2}\mathrm{tr}(K_{A\hat{A}^{\top}}M_{3,1}+K_{\hat{A}A^{\top}}M_{1,3})}{1-\|\sqrt{d}\alpha K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}K_{AA^{\top}}^{\frac{1}{2}}\|_{F}^{2}}$
$=\beta+\frac{\alpha^{2}(1+\mathrm{tr}(K_{AA^{\top}}M^{2}_{1,1}))}{1-\|\sqrt{d}\alpha K_{AA^{\top}}^{\frac{1}{2}}M_{1,1}K_{AA^{\top}}^{\frac{1}{2}}\|_{F}^{2}}$